2.95g of sodium bicarb how much water to add to make 1.0 M solution

4.25g of Sodium carbonate how much water to add to make 1.0 M solution.

Please show all the work very clearly even little steps to help me understand.

"You need to create a 1.0 M solution of sodium bicarbonate (NaHCO3 – baking soda). Your lab kit has a bag of sodium bicarbonate to use for this portion of the experiment.

  1. Write “sodium bicarbonate” on a sample cup and pipette.
  2. Measure a level ½ teaspoon of sodium bicarbonate and place it in the labeled sample cup.
  3. A level ½ teaspoon of sodium bicarbonate has a mass of 2.95g. You will need to determine the volume of distilled water that must be added to this sample in order to create a 1.0M sodium bicarbonate solution. Remember that Molarity = moles of solute/liters of solution. If you are unsure that you are doing your calculation correctly, please check with your instructor before moving on.
  4. Mix the 2.95g of sodium bicarbonate and your calculated volume of distilled water together in the sample cup. Measure the water accurately to the nearest tenth of a milliliter using the graduated cylinder.
  5. Use the pipette to stir until solid is completely dissolved and mixture is clear. This may take several minutes."

thank you
Answer #1

2.95g of sodium bicarbonate how much water to add to make 1.0 M solution

answer

given mass of sodium bicarbonate ( solute) = 2.95g

molar mass of sodium bicarbonate = NaHCO3 = 23+1+12+(16×3) =84

Number of moles = givenmassof solute molarmassof solute =\frac{2.95}{84 } = 0.035

Molarity of the solution = 1.0M

formula to calculate molarity = Numberofmolesof solute volumeofsolutioninliters =

1.0 = 0.035 volumeofsolutioninliters

Volume of solution in liters = 0.035 L

Volume of solution in millilitres =0.035 ×1000= 35mL

therefore 35mL of water is added in 2.95g of NaHCO3 to make a solution of 1.0M

Question : 4.25g of Sodium carbonate how much water to add to make 1.0 M solution.

Answer :

Given mass of sodium Carbonate = 4.25g

Molar mass of sodium Carbonate = (23×2)+12+(16×3)= 106

Number of moles of sodium Carbonate = 0.04 moles

Molarity of the solution = 1.0M

Formula for calculating molarity is Numberofmolesof solute volumeofsolutioninliters

1.0 = 0.04 volumeofsolutioninliters

Volume of solution in liters= 0.04

Volume of solution in millilitres = 0.04× 1000= 40 mL

therefore 40 mL of water is added to 4.25 g of sodium Carbonate to make solution of 1.0 M